# Type 1 Diabetes Management using GLIMMER: Glucose Level Indicator Model with Modified Error Rate

### v1

@@ Abstract
Managing Type 1 Diabetes (T1D) demands constant vigilance as individuals strive to regulate their blood glucose levels to avert the dangers of dysglycemia (hyperglycemia or hypoglycemia). Despite the advent of sophisticated technologies such as automated insulin delivery (AID) systems, achieving optimal glycemic control remains a formidable task. AID systems integrate continuous subcutaneous insulin infusion (CSII) and continuous glucose monitors (CGM) data, offering promise in reducing variability and increasing glucose time-in-range. However, these systems often fail to prevent dysglycemia, partly due to limitations in prediction algorithms that lack the precision to avert abnormal glucose events. This gap highlights the need for proactive behavioral adjustments. We address this need with GLIMMER, Glucose Level Indicator Model with Modified Error Rate, a machine learning approach for forecasting blood glucose levels. GLIMMER categorizes glucose values into normal and abnormal ranges and devises a novel custom loss function to prioritize accuracy in dysglycemic events where patient safety is critical. To evaluate the potential of GLIMMER for T1D management, we both use a publicly available dataset and collect new data involving 25 patients with T1D. In predicting next-hour glucose values, GLIMMER achieved a root mean square error (RMSE) of 23.97 (±3.77) and a mean absolute error (MAE) of 15.83 (±2.09) mg/dL. These results reflect a 23% improvement in RMSE and a 31% improvement in MAE compared to the best-reported error rates.
Index Terms:
@@
Glucose level forecasting, automated insulin delivery, type 1 diabetes, machine learning.I Introduction
@@
Type 1 diabetes (T1D) is an autoimmune condition characterized by the destruction of insulin-producing beta cells in the pancreas, leading to a lifelong dependency on exogenous insulin. Managing T1D is particularly challenging due to the need for continuous monitoring and precise insulin dosing to maintain blood glucose levels within a target range. Poor glucose control can lead to severe complications, including cardiovascular diseases, neuropathy, retinopathy, and kidney failure [1, 2, 3]. Effective management is crucial for enhancing quality of life and reducing the long-term complications associated with T1D. It is estimated that around 8.4 million people worldwide have T1D, representing approximately 5-10% of all diabetes cases [4]. The prevalence of T1D varies significantly across different populations and regions, highlighting the importance of tailored management strategies to address the unique needs of patients globally. Over the years, several advanced technologies have emerged to aid in managing T1D [4]. These technologies aim to improve glycemic control, reduce the risk of complications, and alleviate the daily burden of diabetes management on patients. Each of the following systems has distinct features and functionalities that contribute to more effective diabetes management, addressing different aspects of glucose monitoring and insulin delivery [1, 2, 3].
I-A Type 1 Diabetes Management Technologies
I-A1 Self-Monitoring of Blood Glucose
Self-monitoring of Blood Glucose (SMBG) is essential in managing T1D, allowing patients to regularly measure their blood glucose levels. This enables informed adjustments to diet, insulin therapy, and exercise, helping maintain optimal glucose control and prevent dysglycemia [5, 6]. However, SMBG provides only intermittent data, potentially missing significant glucose fluctuations. Additionally, excessive testing can lead to insulin stacking, increasing the risk of iatrogenic hypoglycemia. Proper patient education and adherence to monitoring schedules are crucial to maximize SMBG benefits while minimizing risks [7, 8].
I-A2 Continuous Glucose Monitoring
Continuous Glucose Monitoring (CGM) systems offer real-time, continuous data on glucose levels by measuring concentrations in the interstitial fluid, closely aligning with plasma glucose values [9]. These systems include a sensor, transmitter, and receiver, providing frequent measurements that enhance glycemic control and enable timely interventions. Most FDA-approved CGM devices allow for non-adjunctive use, enabling therapeutic decisions without additional SMBG verification [10]. Despite their advantages, CGMs have a physiological delay of about 5-6 minutes between blood and interstitial glucose, necessitating occasional SMBG confirmation during rapid glucose changes [11]. In addition, patients must take this delay into account when making decisions about dosing insulin.
I-A3 Hybrid Closed-loop Systems
Hybrid Closed-loop (HCL) systems integrate CGM data with automated insulin delivery, using algorithms to adjust insulin dosing in real-time based on continuous glucose readings [12]. This integration reduces the manual burden on patients and helps maintain glucose levels within the target range more effectively. Clinical trials have shown that HCL systems improve glycemic outcomes, such as lowering HbA1c levels and increasing time in range, while enhancing patients’ quality of life by minimizing the need for constant monitoring and insulin adjustments [13]. As technology advances, HCL systems are expected to offer greater personalization and move closer to fully automated diabetes management solutions.
I-A4 Automated Insulin Delivery Systems
Automated Insulin Delivery (AID) systems combine continuous subcutaneous insulin infusion (CSII) with CGM to dynamically regulate insulin delivery based on real-time glucose data [14, 15, 16]. Utilizing advanced algorithms, AID systems adjust basal insulin rates and administer corrective boluses to maintain glucose levels within desired ranges, reducing the risk of dysglycemia. Systems like the Tandem t:slim X2 with Control-IQ technology111https://www.tandemdiabetes.com/ automate insulin delivery based on sensor data, significantly increasing time in range and decreasing extreme glucose events [17, 18]. Meta-analyses have demonstrated that AID systems improve glucose time in range (TIR) by substantial margins compared to conventional therapy, offering a more user-friendly and effective solution for daily diabetes management [19].
I-B Limitations of Prior Research
Despite advancements in diabetes management technologies, several limitations persist. AID systems often struggle to predict and react quickly to post-meal glucose spikes, delivering corrective boluses that are too slow or insufficient [20, 21]. While CGM systems provide real-time glucose data, they lack predictive features to warn users of impending dysglycemia [22]. Both AID and HCL systems, although integrating CGM and insulin pumps, still require manual inputs for calibration and meals, limiting automation [23, 13]. SMBG, with intermittent testing, can miss critical glucose fluctuations, increasing the risk of undetected dysglycemia [6]. Additionally, the limited availability of clinical data and the high cost of collecting historical data for training create a cold-start issue for researchers.
I-C Proposed Solution
To address these limitations, we propose GLIMMER (Glucose Level Indicator Model with Modified Error Rate), an innovative machine learning model for continuous blood glucose forecasting. GLIMMER uses CGM data, insulin dosages, and meal inputs as key features and significantly improves prediction accuracy through a custom loss function that applies higher penalties in dysglycemic regions, effectively reducing errors in these critical areas and lowering overall prediction penalty to forecasting outcomes that are closely representing abnormal glucose events. By predicting glucose trends in advance, GLIMMER enhances AID systems, enabling proactive self-management behaviors and insulin adjustments. This allows AID systems to modify insulin delivery before glucose levels reach dangerous thresholds, mitigating physiological delays in glucose sensing and insulin action. Additionally, it allows patients to react to the prediction outcomes and actively engage in appropriate dietary, exercise, and insulin injection behaviors to prevent abnormal blood glucose events. As a result, GLIMMER creates opportunities to improve glycemic control and reduce the risk of adverse events, enhancing the safety and effectiveness of AID systems. In addition to developing an innovative glucose forecasting method, we collected approximately 26,707 hours of data from 25 patients with T1D using AID systems. This newly created dataset enables the evaluation of our model in real-world scenarios and addresses the prevalent challenge of limited data availability for experimentation. An example of GLIMMER’s practical application and its potential implementation in clinical practice is illustrated in Fig. 1.
II Related Work
@@
Advancements in forecasting future glucose levels have been crucial for managing patients with T1D, enabling them to proactively respond to glucose fluctuations and significantly improving glucose control. Marigliano et al. [24] demonstrated that integrating predictive alarms with CGM technology reduced hypoglycemic events in adolescents by 40% and severe hypoglycemia by 60%, highlighting the tangible benefits of predictive alerts in real-world settings. Vettoretti et al. [25] further supported these findings by showcasing how artificial intelligence-based diabetes management systems can enhance patient outcomes through early glucose predictions, prompting timely interventions such as insulin dose adjustments or dietary changes to maintain stable glucose levels and mitigate risks like cardiovascular disease and nerve damage. Additionally, Shroff et al. [26] emphasized the shift from reactive to proactive diabetes care with personalized prediction systems that learn individual response patterns, offering tailored alerts to meet each patient’s unique physiological needs. These predictive technologies not only set a new standard in diabetes care by focusing on prevention over treatment but also significantly reduce the daily management burden. Arefeen et al. [27] suggest that machine learning algorithms can effectively predict hyperglycemia events using data from controlled feeding trials. In the following sections, we categorize these algorithms based on their architecture and discuss their approaches to predicting abnormal glucose levels, providing early warnings for timely interventions.
II-A Evidential Deep Learning and Meta-Learning
Machine learning techniques, particularly deep learning algorithms, have achieved reliable glucose-level prediction performance with minimal feature engineering required. Zhu et al. [28] demonstrated the use of evidential deep learning combined with meta-learning to create a model that adapts to individual patient data. This approach significantly enhances prediction accuracy by considering the uncertainty in predictions and personalizing the model to each patient’s unique glucose response patterns. This method’s strength lies in its ability to provide precise predictions with fewer input features, simplifying the data collection process for patients.
II-B Convolutional Recurrent Neural Networks
@@
Another innovative approach is the use of Convolutional Recurrent Neural Networks (CRNN) [29] to estimate glucose levels for up to a 60-minute prediction horizon (PH) based on prior CGM data and information on meal and insulin intakes. Li et al. [30] introduced this model, which combines the feature extraction capabilities of convolutional neural networks (CNN) [31] with the temporal learning capabilities of recurrent neural networks (RNN) [32]. The CRNN model demonstrated superior performance in both simulated and real patient data, providing accurate short-term glucose predictions that are essential for proactive diabetes management.
II-C Long Short-Term Memory Networks
@@
Long Short-Term Memory (LSTM) [33] units have also been employed by Aliberti et al. [34] to predict glucose levels. In this study, they developed a predictive model for blood glucose levels using a multi-patient dataset, focusing on leveraging the strengths of LSTM networks. The researchers compared the performance of LSTM networks with other models like Non-Linear Autoregressive (NAR) neural networks [35] and found that the LSTM model significantly outperformed others in both short- and long-term predictions. The LSTM model demonstrated superior accuracy due to its ability to handle long-term dependencies and mitigate issues like the vanishing gradient problem that commonly affects traditional RNNs. This study’s findings underscore the potential of LSTM networks in enhancing predictive accuracy and clinical outcomes for diabetes management.
II-D Bi-Directional LSTM Variants
Further extending the capabilities of LSTM networks, researchers have explored bi-directional LSTM variants for glucose prediction [36]. Butt et al. [37] investigated how feature transformation techniques could enhance the efficiency of blood glucose prediction models. By employing bi-directional LSTM units, the model can consider both past and future data points, providing a more comprehensive understanding of glucose trends and improving prediction accuracy.
II-E Linear Regression
The Tandem t:slim X2 with Control-IQ technology employs simple linear regression to forecast blood glucose levels 30 minutes ahead based on previous CGM data, highlighting the importance of understanding these algorithms and their accuracy. To address the challenges of applying linear regression to time-series data and multi-step predictions, Zhang et al. [38] developed a multiple linear regression (MLR) model that predicts each future time step separately. This MLR approach combines individual linear regression models, denoted as , each trained to relate the training data to the CGM values at future time points for to . For instance, setting or corresponds to predicting 30 or 60 minutes into the future, respectively, as illustrated in Fig. 2. During the prediction phase, the trained models utilize their respective coefficients and intercepts to forecast glucose levels for the next time steps by applying the models to the testing data one time step prior to the first test point. This iterative process is repeated for each row of the testing matrix, enabling the generation of multi-horizon predictions. Despite the inherent difficulties of using linear regression for time-series forecasting, the structured MLR approach enhances the accuracy of glucose level predictions at various future points, thereby improving the reliability of automated insulin delivery systems. Additionally, this method allows for scalability and adaptability in different clinical settings, making it a valuable tool in the ongoing efforts to optimize diabetes management.
III Methodology
III-A Model Architecture
While many machine learning algorithms are used for event forecasting, CNN-LSTM models excel in continuous multi-modal data by integrating spatially distributed data and capturing time-series patterns. These models are commonly applied in areas such as stock price forecasting, household load prediction, and wind power estimation [39, 40, 41]. Recent studies, including Jaloli and Cescon (2023) [42], have demonstrated that a CNN-LSTM model achieved a lower RMSE for glucose level predictions over a longer forecast horizon compared to other methods.
The decision to use a CNN-LSTM model for GLIMMER is rooted in its hybrid architecture, which combines the strengths of CNNs and LSTMs [31, 33]. The CNN component excels at automatic feature extraction, capturing spatial relationships within the data, while the LSTM component is adept at learning temporal sequences and long-term dependencies. This combination allows the CNN-LSTM model to effectively extract hidden features and correlations among various physiological variables, making it well-suited for forecasting future blood glucose values.
In comparative analyses, the CNN-LSTM model demonstrated superior performance over LSTM, CRNN, and other models, both in terms of predictive accuracy and clinical acceptability [42]. This improved performance is attributed to the model’s sophisticated architecture, which includes stacks of convolutional and LSTM layers capable of learning complex, hidden features in multivariate datasets. The model’s ability to capture rapid and abrupt changes in continuous glucose monitoring trends, due to its capacity to learn intricate dynamics and correlations between variables, further underscores its suitability for this task. However, it is important to note that the effectiveness of the CNN-LSTM model is contingent on the availability of a sufficiently large dataset, which also increases the computational cost compared to simpler reference models. To address the common issue of overfitting in LSTM networks, we incorporate dropout layers after each convolutional or LSTM layer, which has proven effective in enhancing model robustness [43].
III-B Custom Loss Function
Glucose values can be categorized into normal and critical regions, with the latter including hyperglycemia and hypoglycemia thresholds. For patients with T1D, incorrect predictions in these critical regions can be particularly dangerous, resulting in missed intervention opportunities and potentially severe health complications. Despite extensive prior research on predicting glucose levels, a recent review by Woldaregay et al. [44] highlights a significant gap in the field. There is a noticeable lack of comprehensive analysis and modeling concerning the penalties for prediction errors across various dysglycemic regions. Therefore, we propose to penalize the prediction model for errors that occur in critical regions more heavily than those in normal regions. Assuming two thresholds and representing hypoglycemic threshold and hyperglycemia thresholds, respectively, we classify the blood glucose data into three regions: 1 (hypoglycemia) for values below , 2 (normal) for values between and , and 3 (hyperglycemia) for values above . Commonly, insulin delivery devices and research articles set these thresholds to 70 mg/dL for hypoglycemia and 180 mg/dL for hyperglycemia [17]. Although these values may vary slightly from person to person, we use the general values commonly cited in similar studies [24, 25], leaving the in-depth analysis to determine optimal thresholds for each patient for future research. Fig. 3 is an example of CGM readings data that highlights the regions and thresholds.

### v2

@@ Abstract
Managing Type 1 Diabetes (T1D) demands constant vigilance as individuals strive to regulate their blood glucose levels to avert the negative consequences of dysglycemia (hyperglycemia or hypoglycemia). Despite the advent of sophisticated technologies such as automated insulin delivery (AID) systems, achieving optimal glycemic control remains a formidable task. AID systems integrate continuous subcutaneous insulin infusion and continuous glucose monitors (CGM) data, offering promise in reducing variability and improving glucose time-in-range. However, these systems often fail to prevent dysglycemia, partly due to limitations in prediction algorithms that lack the precision to prevent abnormal glucose events. This gap highlights the need for more advanced forecasting methods. We address this need with GLIMMER, Glucose Level Indicator Model with Modified Error Rate, a machine learning approach for forecasting blood glucose levels. GLIMMER categorizes glucose values into normal and abnormal ranges and devises a novel custom loss function to prioritize accuracy in dysglycemic regions where patient safety is critical. To evaluate the potential of GLIMMER for T1D management, we both use a publicly available dataset and collect new data involving 25 patients with T1D. In predicting next-hour glucose values, GLIMMER achieved a root mean square error (RMSE) of 23.97 (3.77) and a mean absolute error (MAE) of 15.83 (2.09) mg/dL. These results reflect a 23% improvement in RMSE and a 31% improvement in MAE compared to the best-reported error rates.
keywords:
@@
Glucose Level Forecasting , Automated Insulin Delivery , Type 1 Diabetes , Machine Learning.[aff1]organization=Arizona State University, addressline=College of Health Solutions, city=Phoenix, postcode=85004, state=AZ, country=USA
[aff2]organization=Arizona State University, addressline=School of Computing and Augmented Intelligence, city=Tempe, postcode=85281, state=AZ, country=USA
[aff3]organization=Mayo Clinic Arizona, addressline=Department of Endocrinology, city=Scottsdale, postcode=85259, state=AZ, country=USA
1 Introduction
@@
Type 1 diabetes (T1D) is an autoimmune condition characterized by the destruction of insulin-producing beta cells in the pancreas, leading to a lifelong dependency on exogenous insulin. Managing T1D is particularly challenging due to the need for continuous monitoring and precise insulin dosing to maintain blood glucose levels within a target range. Poor glucose control can lead to severe complications, including cardiovascular diseases, neuropathy, retinopathy, and kidney failure [1, 2, 3]. As a result, effective diabetes management is crucial for enhancing quality of life and reducing the long-term complications associated with T1D. Furthermore, it is estimated that around 8.4 million people worldwide suffer from T1D, representing approximately 5-10% of all diabetes cases [4]. The prevalence of T1D varies significantly across different populations and regions, highlighting the importance of tailored management strategies to address the unique needs of patients globally.
Predicting future blood glucose levels is crucial for preventing dangerous dysglycemic events (hypoglycemia and hyperglycemia) and supporting timely, proactive interventions focused on medication dosing and behavioral modifications. Errors in these critical regions carry serious health risks, whereas deviations within the normal range are generally less impactful [5, 6]. Long-term predictions, such as those extending beyond the more common 15–30 minute horizon to 60 minutes or longer, offer patients and clinicians greater lead time to respond effectively, enhancing safety and reducing stress [7, 8, 9]. When embedded in technologies like automated insulin delivery (AID) and hybrid closed-loop (HCL) systems, robust prediction models can help overcome the physiological delays of insulin action and glucose sensing, allowing for earlier alerts or therapeutic adjustments before glucose levels enter dangerous ranges [10, 11, 12].
Despite advancements in diabetes management technologies, several limitations persist. Existing glucose prediction models often rely on generic architectures and overlook the domain-specific complexities of 1 T1D management. Many models fail to maintain accuracy in clinically significant dysglycemic regions, particularly during rapid fluctuations such as postprandial glucose spikes, where timely intervention is critical. AID systems, which rely heavily on these forecasts, often struggle to predict and react quickly enough to post-meal rises, leading to corrective boluses that are delayed or insufficient [13, 14]. Another key challenge lies in data availability: the high cost and burden of collecting detailed, labeled data limit access to large, high-quality datasets [15]. This creates a cold-start problem for researchers and hinders model generalizability and personalization [16].
To address these challenges, we propose GLIMMER111https://github.com/SamanKhamesian/GLIMMER (Glucose Level Indicator Model with Modified Error Rate), a domain-specific prediction framework designed for T1D management. Our approach uniquely combines tailored feature engineering, extensive hyperparameter tuning, and a novel custom loss function that prioritizes accuracy in clinically critical (dysglycemic) regions—an objective not addressed by prior studies. This loss function places greater weight on errors during hyperglycemic and hypoglycemic events, enabling the model to more effectively anticipate postprandial glucose spikes and support earlier corrective actions. These improvements directly address a major limitation of current AID systems, which often respond too slowly or deliver insufficient insulin after meals [13, 14].
To overcome the limitations of existing datasets, we also conducted a real-world study in collaboration with Mayo Clinic Arizona and introduced the AZT1D [17]222https://data.mendeley.com/datasets/gk9m674wcx/1, a dataset comprising 26,707 hours of multimodal data from 25 patients with T1D. GLIMMER demonstrates superior forecasting performance even at extended prediction horizons (60 minutes), achieving a 23% improvement in RMSE and a 31% improvement in MAE on the OhioT1DM dataset, and generalizes effectively to the AZT1D dataset. We illustrate its potential integration into clinical workflows in Figure 1.
2 Related Work
@@
Machine learning and deep learning approaches have significantly advanced blood glucose prediction, enabling proactive interventions and improved glycemic control for patients with T1D. Marigliano et al. [8] demonstrated that integrating predictive alarms with CGM technology reduced hypoglycemic events in adolescents by 40% and severe hypoglycemia by 60%, highlighting the clinical benefits of early warnings. Similarly, Vettoretti et al. [9] showcased how artificial intelligence-based management systems can improve outcomes by prompting timely interventions such as insulin dose adjustments or dietary changes. Building on this idea, Shroff et al. [18] proposed a personalized prediction system that learns individual response patterns and provides tailored alerts, shifting diabetes care from reactive to proactive. In line with these efforts, Arefeen et al. [19] developed machine learning algorithms to predict postprandial hyperglycemia using data from controlled feeding trials.
@@
To further improve prediction capabilities, researchers have explored more sophisticated architectures such as Convolutional Recurrent Neural Networks (CRNN) [20], which estimate glucose levels for up to a 60-minute prediction horizon (PH) based on prior CGM data and information on meal and insulin intakes. Li et al. [21] introduced this model, combining the feature extraction capabilities of convolutional neural networks (CNN) [22] with the temporal learning capabilities of recurrent neural networks (RNN) [23]. The CRNN demonstrated superior performance in both simulated and real patient data, providing accurate short-term glucose predictions essential for proactive diabetes management.
@@
Along similar lines, Long Short-Term Memory (LSTM) [24] networks have also been employed to capture long-term dependencies in glucose trends. Aliberti et al. [25] developed a predictive model using a multi-patient dataset and demonstrated that LSTM networks significantly outperformed Non-Linear Autoregressive (NAR) neural networks [26] in both short- and long-term predictions. This improvement is attributed to LSTM’s ability to handle long-term dependencies and mitigate issues like the vanishing gradient problem, underscoring its potential to enhance predictive accuracy and clinical outcomes for diabetes management. Building upon LSTM networks, bi-directional variants have been explored to further improve predictions by considering both past and future data points. Sun et al. [27] applied bi-directional LSTM networks for glucose prediction, while Butt et al. [28] investigated how feature transformation techniques can enhance the efficiency of such models. These approaches provide a more comprehensive understanding of glucose trends and lead to improved prediction accuracy.
In contrast to deep learning methods, linear regression remains a simple yet effective approach in commercial systems. The Tandem t:slim X2 with Control-IQ technology, for example, employs linear regression to forecast glucose levels 30 minutes ahead based on previous CGM data. To address the limitations of linear regression in time-series and multi-step predictions, Zhang et al. [29] developed a multiple linear regression (MLR) model that predicts each future time step separately. Despite the challenges of using linear regression for time-series forecasting, the structured MLR approach improves prediction accuracy at various horizons, thereby enhancing the reliability of automated insulin delivery systems.
While these studies have considerably advanced prediction accuracy and clinical utility, many rely on generic architectures without clear justification of design choices or incorporation of domain-specific features such as meal timing or physical activity. Furthermore, most models optimize for mean prediction error and overlook the clinical importance of minimizing errors in hypo- and hyperglycemic regions [6, 8]. A summary of these models, their main characteristics, and their limitations is presented in Table 1, illustrating the need for a domain-specific, clinically-focused approach.
| Model | Key Features | Limitations |
| FCNN [16] | Fully-connected neural network; simple, fast, easy to train. | Lacks the ability to capture temporal dependencies important for glucose trends. |
| CNN [30] | Local pattern extraction via convolutions; low latency. | Limited in modeling long-term dependencies; emphasizes mean error rather than clinically relevant regions. |
| CRNN [21] | Combines CNN and RNN to capture local & temporal patterns. | Limited attention to dysglycemic regions; general-purpose architecture without domain-specific enhancements. |
| Bi-LSTM [27] | Uses past and future context for improved sequence modeling. | Computationally intensive; does not incorporate domain-specific loss optimization. |
| GRU [31] | Faster and lighter than LSTM; good for medium-range dependencies. | Less effective at modeling long sequences; prioritizes mean error over critical glucose regions. |
| Transformer [32] | Attention-based; models long-range dependencies; handles multimodal inputs. | Requires large datasets; does not explicitly focus on hypo- and hyperglycemic regions. |
| ARIMA [33] | Linear autoregressive; interpretable baseline. | Limited in modeling nonlinear, multimodal, and long-term dynamics. |
| MLP [34] | Simple shallow network; fast inference of basic nonlinear patterns. | Lacks capacity to model sequential dependencies and specific glucose dynamics. |
3 Methodology
3.1 Model Architecture
The decision to use a CNN-LSTM model for GLIMMER is motivated by its hybrid architecture, which combines the automatic feature extraction capabilities of convolutional layers with the temporal modeling strengths of LSTM units [22, 24]. CNNs capture spatial relationships within multivariate physiological data, while LSTMs are well-suited for learning sequential patterns and long-term dependencies. This integration enables the model to uncover hidden features and capture abrupt changes in CGM trends, making it highly effective for blood glucose forecasting. Recent work by Jaloli and Cescon (2023) [7] demonstrated that CNN-LSTM outperforms LSTM, CRNN, and other models in both predictive accuracy and clinical relevance, especially over longer prediction horizons. Similar success has been observed in other time-series domains such as stock forecasting, energy load prediction, and wind power estimation [35, 36, 37]. To find the most suitable parameters for the CNN-LSTM architecture, we designed an experiment to analyze and identify the optimal configuration, with the fine-tuning process detailed in Section 5.1.
3.2 Custom Loss Function
Glucose values are typically divided into three clinically meaningful regions: hypoglycemia, normal, and hyperglycemia as shown in (1). Values below the hypoglycemia threshold indicate dangerously low glucose levels, while values above the hyperglycemia threshold indicate excessively high levels. Measurements that fall between these two thresholds are considered within the normal range. In practice, and are often set to 70 mg/dL and 180 mg/dL, respectively, based on guidelines and prior studies [38, 8, 9]. Prediction errors in the critical regions of hypo- and hyperglycemia are particularly dangerous for patients with T1D, as they can lead to missed interventions and serious complications. Woldaregay et al. [5] highlighted that prior studies rarely account for the uneven clinical impact of such errors.
To address this gap, we propose penalizing prediction errors more heavily in critical regions than in the normal range. Figure 2 is an example of CGM readings data that highlights the regions and thresholds.